# SURE-VQA: Systematic Understanding of Robustness Evaluation in Medical VQA Tasks

### v1

@@ C.3.4 Ratio of unique questions in the datasets
C.4 Detailed Results of the Robustness Study
@@
Tables 14-19 show the detailed results of the robustness study. Further, Figure 12 shows the inter-method and inter-shift variability of the different PEFT methods, so not including full FT.
@@
No Finetune 0.59 0.29 0.49 2.91 2.79 0.96 0.54 0.47 0.87 2.86 3.16 1.11
Full Finetune 0.57 0.48 0.83 4.03 3.17 0.79 0.56 0.35 0.63 4.11 4.38 1.07
Prompt 0.85$\pm$0.01 0.62$\pm$0.03 0.73$\pm$0.05 4.22$\pm$0.04 3.69$\pm$0.06 0.87$\pm$0.02 0.85$\pm$0.01 0.49$\pm$0.12 0.57$\pm$0.14 4.17$\pm$0.01 4.35$\pm$0.16 1.04$\pm$0.04
LoRA 0.88$\pm$0.0 0.45$\pm$0.04 0.51$\pm$0.04 4.34$\pm$0.06 3.61$\pm$0.06 0.83$\pm$0.03 0.87$\pm$0.0 0.39$\pm$0.07 0.45$\pm$0.08 4.26$\pm$0.01 4.26$\pm$0.07 1.0$\pm$0.02
$\mathtt{(IA)^{3}}$ 0.85$\pm$0.01 0.64$\pm$0.07 0.75$\pm$0.08 4.35$\pm$0.02 3.4$\pm$0.15 0.78$\pm$0.04 0.87$\pm$0.02 0.53$\pm$0.06 0.61$\pm$0.06 4.23$\pm$0.01 4.21$\pm$0.27 0.99$\pm$0.07
Most Freq. 0.69 - - 3.22 - - 0.696 - - 3.05 - -
@@
No Finetune 0.46 0.35 0.75 2.13 2.33 1.1 0.46 0.29 0.64 2.22 2.03 0.91
Prompt 0.55$\pm$0.01 0.5$\pm$0.01 0.91$\pm$0.03 3.24$\pm$0.0 1.88$\pm$0.06 0.58$\pm$0.02 0.59$\pm$0.05 0.55$\pm$0.07 0.94$\pm$0.18 3.18$\pm$0.04 2.26$\pm$0.6 0.71$\pm$0.19
LoRA 0.64$\pm$0.03 0.47$\pm$0.07 0.73$\pm$0.11 3.24$\pm$0.02 1.93$\pm$0.05 0.6$\pm$0.02 0.69$\pm$0.01 0.53$\pm$0.18 0.76$\pm$0.25 3.12$\pm$0.03 2.95$\pm$0.0 0.94$\pm$0.01
$\mathtt{(IA)^{3}}$ 0.55$\pm$0.01 0.47$\pm$0.02 0.85$\pm$0.03 3.26$\pm$0.01 1.87$\pm$0.02 0.57$\pm$0.01 0.55$\pm$0.08 0.53$\pm$0.06 0.96$\pm$0.09 3.15$\pm$0.02 2.64$\pm$0.53 0.84$\pm$0.17
@@
No Finetune 0.42 0.4 0.96 2.4 2.45 1.02 0.43 0.33 0.75 2.39 1.94 0.81
Full Finetune 0.7 0.55 0.77 3.16 2.23 0.71 0.76 0.08 0.1 2.97 1.02 0.34
Prompt 0.86$\pm$0.0 0.75$\pm$0.01 0.87$\pm$0.01 3.12$\pm$0.02 2.38$\pm$0.02 0.76$\pm$0.01 0.82$\pm$0.04 0.86$\pm$0.01 1.05$\pm$0.06 2.9$\pm$0.01 1.7$\pm$0.02 0.59$\pm$0.01
LoRA 0.86$\pm$0.01 0.77$\pm$0.0 0.9$\pm$0.01 3.23$\pm$0.02 2.47$\pm$0.05 0.76$\pm$0.02 0.84$\pm$0.0 0.74$\pm$0.06 0.88$\pm$0.07 3.09$\pm$0.03 1.72$\pm$0.11 0.56$\pm$0.04
$\mathtt{(IA)^{3}}$ 0.83$\pm$0.02 0.75$\pm$0.01 0.9$\pm$0.02 3.21$\pm$0.05 2.46$\pm$0.04 0.77$\pm$0.0 0.8$\pm$0.02 0.72$\pm$0.11 0.91$\pm$0.15 2.98$\pm$0.06 1.51$\pm$0.05 0.51$\pm$0.02
Most Freq. 0.75 - - 2.57 - - 0.73 - - 2.23 - -
@@
No Finetune 0.42 0.36 0.85 1.37 2.02 1.47 0.41 0.4 0.97 1.39 1.29 0.93
Prompt 0.74$\pm$0.01 0.69$\pm$0.02 0.93$\pm$0.01 2.63$\pm$0.1 2.12$\pm$0.07 0.81$\pm$0.01 0.67$\pm$0.03 0.44$\pm$0.01 0.66$\pm$0.02 2.35$\pm$0.06 1.26$\pm$0.07 0.54$\pm$0.02
LoRA 0.74$\pm$0.0 0.7$\pm$0.0 0.95$\pm$0.01 2.67$\pm$0.16 2.14$\pm$0.01 0.81$\pm$0.05 0.73$\pm$0.0 0.43$\pm$0.03 0.6$\pm$0.04 2.36$\pm$0.02 1.29$\pm$0.09 0.55$\pm$0.04
$\mathtt{(IA)^{3}}$ 0.74$\pm$0.0 0.69$\pm$0.02 0.93$\pm$0.02 2.74$\pm$0.07 2.13$\pm$0.04 0.78$\pm$0.03 0.7$\pm$0.04 0.39$\pm$0.06 0.56$\pm$0.1 2.36$\pm$0.02 1.28$\pm$0.06 0.54$\pm$0.02

### v2

@@ C.1.4 Ratio of unique questions in the datasets
C.2 Hyperparameter Search
We performed several hyperparameter sweeps for each dataset and PEFT method in order to find suitable setups for the experiments in the FT robustness study. For the hyperparameter sweeps, we trained on the whole training set for each dataset and PEFT method and ran inference on the validation set. Training ran for 3 epochs and 3 seeds for each experiment.
C.2.1 Prompt Tuning
  • $\bullet$
    Number of tokens: $[40,60,80,100]$
  • $\bullet$
    Learning rate: $[3e-2,3e-1]$
Table 5: Hyperparameter sweep for prompt tuning on the SLAKE dataset. Selected hyperparameters for the final FT robustness study are highlighted. Mean and standard deviation are reported for three seeds.
(a) Hyperparameter sweep for the medical base model.
# Tokens Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
40 3e-2 0.79 +/- 0.02 4.21 +/- 0.03
40 3e-1 0.8 +/- 0.02 4.24 +/- 0.06
60 3e-2 0.76 +/- 0.04 4.22 +/- 0.03
60 3e-1 0.81 +/- 0.01 4.22 +/- 0.06
80 3e-2 0.78 +/- 0.05 4.21 +/- 0.02
80 3e-1 0.82 +/- 0.01 4.23 +/- 0.02
100 3e-2 0.77 +/- 0.06 4.21 +/- 0.05
100 3e-1 0.81 +/- 0.0 4.23 +/- 0.04
(b) Hyperparameter sweep for the non-medical base model.
# Tokens Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
40 3e-2 0.76 +/- 0.04 4.15 +/- 0.02
40 3e-1 0.79 +/- 0.01 4.18 +/- 0.02
60 3e-2 0.77 +/- 0.0 4.13 +/- 0.02
60 3e-1 0.79 +/- 0.01 4.17 +/- 0.02
80 3e-2 0.76 +/- 0.01 4.14 +/- 0.05
80 3e-1 0.78 +/- 0.01 4.17 +/- 0.01
100 3e-2 0.77 +/- 0.02 4.17 +/- 0.02
100 3e-1 0.79 +/- 0.02 4.17 +/- 0.06
Table 6: Hyperparameter sweep for prompt tuning on the OVQA dataset. Selected hyperparameters for the final FT robustness study are highlighted. Mean and standard deviation are reported for three seeds.
(a) Hyperparameter sweep for the medical base model.
# Tokens Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
40 3e-2 0.85 +/- 0.0 3.05 +/- 0.05
40 3e-1 0.85 +/- 0.01 3.05 +/- 0.04
60 3e-2 0.85 +/- 0.01 3.06 +/- 0.03
60 3e-1 0.85 +/- 0.0 3.05 +/- 0.04
80 3e-2 0.82 +/- 0.03 3.05 +/- 0.05
80 3e-1 0.84 +/- 0.01 3.07 +/- 0.03
100 3e-2 0.83 +/- 0.02 3.1 +/- 0.03
100 3e-1 0.85 +/- 0.0 3.08 +/- 0.02
(b) Hyperparameter sweep for the non-medical base model.
# Tokens Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
40 3e-2 0.8 +/- 0.03 2.97 +/- 0.02
40 3e-1 0.84 +/- 0.01 2.98 +/- 0.03
60 3e-2 0.81 +/- 0.02 2.97 +/- 0.06
60 3e-1 0.83 +/- 0.0 3.01 +/- 0.05
80 3e-2 0.81 +/- 0.03 3.04 +/- 0.05
80 3e-1 0.84 +/- 0.0 3.0 +/- 0.05
100 3e-2 0.81 +/- 0.01 3.0 +/- 0.06
100 3e-1 0.84 +/- 0.01 3.03 +/- 0.07
Table 7: Hyperparameter sweep for prompt tuning on the MIMIC dataset. Selected hyperparameters for the final FT robustness study are highlighted. Mean and standard deviation are reported for three seeds.
(a) Hyperparameter sweep for the medical base model.
# Tokens Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
40 3e-2 0.67 +/- 0.02 3.21 +/- 0.03
40 3e-1 0.67 +/- 0.01 3.2 +/- 0.05
60 3e-2 0.68 +/- 0.01 3.18 +/- 0.01
60 3e-1 0.69 +/- 0.01 3.24 +/- 0.02
80 3e-2 0.66 +/- 0.05 3.21 +/- 0.02
80 3e-1 0.69 +/- 0.02 3.21 +/- 0.03
100 3e-2 0.68 +/- 0.02 3.23 +/- 0.03
100 3e-1 0.67 +/- 0.01 3.21 +/- 0.03
(b) Hyperparameter sweep for the non-medical base model.
# Tokens Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
40 3e-2 0.69 +/- 0.0 3.16 +/- 0.03
40 3e-1 0.69 +/- 0.0 3.18 +/- 0.03
60 3e-2 0.69 +/- 0.0 3.22 +/- 0.04
60 3e-1 0.68 +/- 0.03 3.19 +/- 0.02
80 3e-2 0.68 +/- 0.01 3.2 +/- 0.01
80 3e-1 0.7 +/- 0.0 3.19 +/- 0.04
100 3e-2 0.69 +/- 0.01 3.21 +/- 0.04
100 3e-1 0.68 +/- 0.01 3.17 +/- 0.02
C.2.2 LoRA
For LoRA, we performed the following hyperparameter sweeps:
  • $\bullet$
    Rank: $[16,32,64,128,256]$
  • $\bullet$
    Learning rate: $[3e-5,3e-4]$
$\alpha$ is set to $2\times\text{Rank}$. The results for SLAKE can be found in Table 8, for OVQA in Table 9, and for MIMIC in Table 10. Note that some of the hyperparameter configurations led to instabilities during training loss, indicated by "NaN".
Table 8: Hyperparameter sweep for LoRA on the SLAKE dataset. Selected hyperparameters for the final FT robustness study are highlighted. Mean and standard deviation are reported for three seeds. Rows with "NaN" showed instabilities in the loss during training.
(a) Hyperparameter sweep for the medical base model.
Rank Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
16 3e-5 0.83 +/- 0.01 4.29 +/- 0.02
16 3e-4 0.82 +/- 0.01 4.27 +/- 0.04
32 3e-5 0.85 +/- 0.01 4.31 +/- 0.03
32 3e-4 0.73 +/- 0.07 4.25 +/- 0.01
64 3e-5 0.84 +/- 0.01 4.33 +/- 0.04
64 3e-4 0.51 +/- 0.07 2.92 +/- 1.41
128 3e-5 0.84 +/- 0.01 4.34 +/- 0.02
128 3e-4 NaN NaN
256 3e-5 0.83 +/- 0.01 4.31 +/- 0.01
256 3e-4 NaN NaN
(b) Hyperparameter sweep for the non-medical base model.
Rank Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
16 3e-5 0.84 +/- 0.01 4.25 +/- 0.02
16 3e-4 0.82 +/- 0.01 4.26 +/- 0.02
32 3e-5 0.84 +/- 0.0 4.33 +/- 0.02
32 3e-4 0.78 +/- 0.01 4.2 +/- 0.05
64 3e-5 0.86 +/- 0.01 4.34 +/- 0.02
64 3e-4 0.45 +/- 0.09 1.86 +/- 0.78
128 3e-5 0.86 +/- 0.01 4.32 +/- 0.04
128 3e-4 NaN NaN
256 3e-5 0.84 +/- 0.01 4.31 +/- 0.03
256 3e-4 NaN NaN
Table 9: Hyperparameter sweep for LoRA on the OVQA dataset. Selected hyperparameters for the final FT robustness study are highlighted. Mean and standard deviation are reported for three seeds. Rows with "NaN" showed instabilities in the loss during training.
(a) Hyperparameter sweep for the medical base model.
Rank Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
16 3e-5 0.84 +/- 0.0 3.13 +/- 0.06
16 3e-4 0.83 +/- 0.02 3.18 +/- 0.02
32 3e-5 0.85 +/- 0.0 3.15 +/- 0.01
32 3e-4 0.82 +/- 0.01 3.08 +/- 0.04
64 3e-5 0.85 +/- 0.0 3.2 +/- 0.02
64 3e-4 0.66 +/- 0.01 2.03 +/- 0.13
128 3e-5 0.85 +/- 0.0 3.18 +/- 0.03
128 3e-4 NaN NaN
256 3e-5 0.85 +/- 0.0 3.18 +/- 0.02
256 3e-4 NaN NaN
(b) Hyperparameter sweep for the non-medical base model.
Rank Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
16 3e-5 0.84 +/- 0.0 3.08 +/- 0.03
16 3e-4 0.81 +/- 0.02 3.17 +/- 0.02
32 3e-5 0.84 +/- 0.01 3.12 +/- 0.04
32 3e-4 0.81 +/- 0.03 3.07 +/- 0.03
64 3e-5 0.84 +/- 0.01 3.15 +/- 0.01
64 3e-4 0.61 +/- 0.07 2.12 +/- 0.11
128 3e-5 0.84 +/- 0.0 3.2 +/- 0.04
128 3e-4 NaN NaN
256 3e-5 0.84 +/- 0.01 3.2 +/- 0.05
256 3e-4 NaN NaN
Table 10: Hyperparameter sweep for LoRA on the MIMIC dataset. Selected hyperparameters for the final FT robustness study are highlighted. Mean and standard deviation are reported for three seeds. Rows with "NaN" showed instabilities in the loss during training.
(a) Hyperparameter sweep for the medical base model.
Rank Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
16 3e-5 0.71 +/- 0.01 3.33 +/- 0.02
16 3e-4 0.68 +/- 0.01 3.23 +/- 0.05
32 3e-5 0.71 +/- 0.0 3.36 +/- 0.01
32 3e-4 0.42 +/- 0.18 2.43 +/- 0.07
64 3e-5 0.71 +/- 0.01 3.35 +/- 0.02
64 3e-4 NaN NaN
128 3e-5 0.71 +/- 0.0 3.37 +/- 0.02
128 3e-4 NaN NaN
256 3e-5 NaN NaN
256 3e-4 NaN NaN
(b) Hyperparameter sweep for the non-medical base model.
Rank Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
16 3e-5 0.7 +/- 0.01 3.28 +/- 0.03
16 3e-4 0.68 +/- 0.01 3.25 +/- 0.05
32 3e-5 0.7 +/- 0.0 3.34 +/- 0.04
32 3e-4 0.44 +/- 0.07 2.43 +/- 0.05
64 3e-5 0.71 +/- 0.01 3.34 +/- 0.02
64 3e-4 NaN NaN
128 3e-5 0.7 +/- 0.0 3.36 +/- 0.03
128 3e-4 NaN NaN
256 3e-5 0.7 +/- 0.02 3.37 +/- 0.02
256 3e-4 NaN NaN
C.3 (IA)3
  • $\bullet$
    Learning rate: $[3e-3,3e-2,3e-1]$
Table 11: Hyperparameter sweep for $\mathtt{(IA)^{3}}$on the SLAKE dataset. Selected hyperparameters for the final FT robustness study are highlighted. Mean and standard deviation are reported for three seeds.
(a) Hyperparameter sweep for the medical base model.
Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
lr3e-3 0.63 +/- 0.02 3.72 +/- 0.01
lr3e-2 0.83 +/- 0.01 4.32 +/- 0.03
lr3e-1 0.65 +/- 0.01 4.24 +/- 0.04
(b) Hyperparameter sweep for the non-medical base model.
Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
lr3e-3 0.75 +/- 0.01 3.61 +/- 0.01
lr3e-2 0.86 +/- 0.01 4.27 +/- 0.02
lr3e-1 0.66 +/- 0.03 4.18 +/- 0.04
Table 12: Hyperparameter sweep for $\mathtt{(IA)^{3}}$on the OVQA dataset. Selected hyperparameters for the final FT robustness study are highlighted. Mean and standard deviation are reported for three seeds.
(a) Hyperparameter sweep for the medical base model.
Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
lr3e-3 0.75 +/- 0.01 2.91 +/- 0.04
lr3e-2 0.84 +/- 0.0 3.15 +/- 0.03
lr3e-1 0.8 +/- 0.04 3.08 +/- 0.04
(b) Hyperparameter sweep for the non-medical base model.
Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
lr3e-3 0.75 +/- 0.02 2.65 +/- 0.03
lr3e-2 0.84 +/- 0.0 3.1 +/- 0.05
lr3e-1 0.8 +/- 0.02 3.05 +/- 0.03
Table 13: Hyperparameter sweep for $\mathtt{(IA)^{3}}$on the MIMIC dataset. Selected hyperparameters for the final FT robustness study are highlighted. Mean and standard deviation are reported for three seeds.
(a) Hyperparameter sweep for the medical base model.
Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
lr3e-3 0.53 +/- 0.01 2.92 +/- 0.01
lr3e-2 0.7 +/- 0.01 3.33 +/- 0.02
lr3e-1 0.61 +/- 0.05 3.1 +/- 0.03
(b) Hyperparameter sweep for the non-medical base model.
Learning Rate Closed-Ended (Gemma Accuracy) Open-Ended (Gemma Score)
lr3e-3 0.6 +/- 0.01 2.7 +/- 0.02
lr3e-2 0.7 +/- 0.0 3.33 +/- 0.02
lr3e-1 0.6 +/- 0.11 3.0 +/- 0.06
C.4 Robustness Results for the Non-Medical Base Model
The robustness study results for the non-medical base model are presented in Figure 16 and Figure 17, corresponding to Figure 7 and Figure 8 for the medical base model. The main trends analyzed in section 3.4 for the medical model also apply to the non-medical base model, as indicated by Figure 6, which suggests similar behavior across the two models.
While the overall trends are consistent, some minor differences are observed. For example, LoRA is no longer an outlier for closed-ended questions in SLAKE. However, a similar trend appears for the non-medical model on OVQA, where LoRA shows a lower RR compared to Prompt and $\mathtt{(IA)^{3}}$, primarily due to the question type shift. Additionally, while the question type shift continues to cause significant failures, this effect is now limited to the OVQA dataset, affecting both open-ended and closed-ended questions. Finally, while the trends for full FT are similar on the SLAKE and the OVQA dataset between the medical and the non-medical base model, the non-medical base model performs worse on the MIMIC dataset than its medical counterpart.
Despite these differences, the results confirm the main insights from the robustness study: None of the methods consistently outperforms others in terms of robustness, and robustness trends are more consistent across FT methods than across shifts. LoRA remains the best-performing PEFT method overall on the i.i.d. data.
Figure 16: Results of the FT Robustness Study on the i.i.d. and OoD Test Set for the Non-Medical Base Model. Reported results show the mean over three seeds (exception: no FT, full FT) with the standard deviation for the non-baselines. Gemma Accuracy refers to the accuracy being rated by Gemma. For MIMIC, the most frequent sanity baseline can not be calculated as too few questions match the training set.
Figure 17: Results of the FT Robustness Study Focusing on the Relative Robustness (RR) for the Non-Medical Base Model. Since the study focuses on comparing the FT method and our definition of OoD holds for these methods, the no FT baseline is excluded. (a) RR on the three datasets for the methods. Results show the mean and standard deviation over three seeds (exception: full FT). (b) RR Ranking of all FT methods. (c) Standard deviation between shifts vs. standard deviation between FT methods.
C.5 Detailed Results of the Robustness Study
@@
Tables 14-19 show the detailed results of the robustness study. Further, Figure 18 shows the inter-method and inter-shift variability of the different PEFT methods for the medical base model, so not including full FT.
@@
No Finetuning 0.6 0.31 0.51 2.57 2.6 1.01 0.54 0.71 1.3 2.53 3.1 1.23
Full Finetuning 0.57 0.48 0.84 4.09 3.25 0.79 0.56 0.35 0.63 4.14 4.38 1.06
Prompt Tuning 0.85 +/- 0.01 0.62 +/- 0.03 0.73 +/- 0.05 4.27 +/- 0.03 3.63 +/- 0.03 0.85 +/- 0.01 0.85 +/- 0.01 0.49 +/- 0.12 0.57 +/- 0.14 4.23 +/- 0.01 4.34 +/- 0.12 1.03 +/- 0.03
LoRA 0.88 +/- 0.0 0.45 +/- 0.04 0.51 +/- 0.04 4.36 +/- 0.06 3.59 +/- 0.07 0.82 +/- 0.03 0.87 +/- 0.0 0.39 +/- 0.07 0.45 +/- 0.08 4.29 +/- 0.02 4.15 +/- 0.07 0.97 +/- 0.01
IA3 0.85 +/- 0.01 0.64 +/- 0.07 0.75 +/- 0.08 4.36 +/- 0.03 3.46 +/- 0.12 0.79 +/- 0.03 0.87 +/- 0.02 0.53 +/- 0.06 0.61 +/- 0.06 4.27 +/- 0.01 4.15 +/- 0.22 0.97 +/- 0.05
Most Frequent 0.69 - - 3.25 - - 0.696 - - 3.1 - -
(b) Results for the non-medical base model.
@@
No Finetuning 0.56 0.75 1.34 2.42 2.38 0.98 0.63 0.41 0.66 2.39 2.67 1.11
Full Finetuning 0.52 0.48 0.91 4.05 3.27 0.81 0.73 0.47 0.64 4.07 4.69 1.15
Prompt Tuning 0.8 +/- 0.01 0.58 +/- 0.04 0.72 +/- 0.06 4.27 +/- 0.05 3.28 +/- 0.05 0.77 +/- 0.02 0.83 +/- 0.01 0.65 +/- 0.06 0.78 +/- 0.06 4.21 +/- 0.05 3.85 +/- 0.2 0.91 +/- 0.04
LoRA 0.88 +/- 0.02 0.68 +/- 0.01 0.77 +/- 0.02 4.37 +/- 0.06 3.28 +/- 0.07 0.75 +/- 0.02 0.86 +/- 0.0 0.69 +/- 0.07 0.8 +/- 0.08 4.25 +/- 0.03 4.61 +/- 0.01 1.08 +/- 0.01
IA3 0.86 +/- 0.02 0.57 +/- 0.08 0.66 +/- 0.08 4.31 +/- 0.05 3.28 +/- 0.02 0.76 +/- 0.01 0.87 +/- 0.01 0.57 +/- 0.09 0.65 +/- 0.09 4.28 +/- 0.02 4.26 +/- 0.13 1.0 +/- 0.03
Most Frequent 0.69 - - 3.25 - - 0.696 - - 3.1 - -
(a) Results for the medical base model.
Modality Shift Question Type Shift
OoD: X-Ray OoD: Size
Closed-Ended Open-Ended Closed-Ended Open-Ended
i.i.d. OoD RR i.i.d. OoD RR i.i.d. OoD RR i.i.d. OoD RR
No Finetuning 0.48 0.35 0.73 1.73 1.85 1.07 0.46 0.59 1.28 1.77 2.03 1.15
Prompt Tuning 0.55 +/- 0.01 0.5 +/- 0.01 0.91 +/- 0.03 3.24 +/- 0.01 1.96 +/- 0.05 0.6 +/- 0.01 0.59 +/- 0.05 0.55 +/- 0.07 0.94 +/- 0.18 3.18 +/- 0.03 2.26 +/- 0.6 0.71 +/- 0.19
LoRA 0.64 +/- 0.03 0.47 +/- 0.07 0.73 +/- 0.11 3.25 +/- 0.03 2.03 +/- 0.04 0.62 +/- 0.02 0.69 +/- 0.01 0.53 +/- 0.18 0.76 +/- 0.25 3.13 +/- 0.02 2.95 +/- 0.0 0.94 +/- 0.01
IA3 0.55 +/- 0.01 0.47 +/- 0.02 0.85 +/- 0.03 3.27 +/- 0.02 1.94 +/- 0.03 0.6 +/- 0.01 0.55 +/- 0.08 0.55 +/- 0.09 0.99 +/- 0.09 3.16 +/- 0.02 2.64 +/- 0.53 0.84 +/- 0.17
(b) Results for the non-medical base model.
Modality Shift Question Type Shift
OoD: X-Ray OoD: Size
Closed-Ended Open-Ended Closed-Ended Open-Ended
i.i.d. OoD RR i.i.d. OoD RR i.i.d. OoD RR i.i.d. OoD RR
No Finetuning 0.43 0.32 0.75 1.5 1.61 1.08 0.4 0.35 0.88 1.53 2.03 1.32
Prompt Tuning 0.64 +/- 0.03 0.65 +/- 0.03 1.01 +/- 0.09 3.27 +/- 0.01 2.04 +/- 0.02 0.62 +/- 0.01 0.65 +/- 0.02 0.61 +/- 0.07 0.94 +/- 0.12 3.17 +/- 0.03 2.95 +/- 0.0 0.93 +/- 0.01
LoRA 0.67 +/- 0.01 0.46 +/- 0.06 0.68 +/- 0.08 3.25 +/- 0.02 2.01 +/- 0.04 0.62 +/- 0.01 0.68 +/- 0.01 0.69 +/- 0.03 1.01 +/- 0.06 3.15 +/- 0.02 2.26 +/- 0.6 0.72 +/- 0.19
IA3 0.67 +/- 0.0 0.65 +/- 0.07 0.97 +/- 0.1 3.3 +/- 0.04 2.02 +/- 0.04 0.61 +/- 0.01 0.68 +/- 0.02 0.57 +/- 0.09 0.83 +/- 0.14 3.17 +/- 0.01 2.33 +/- 0.53 0.74 +/- 0.17
@@
No Finetuning 0.4 0.4 0.99 2.27 2.27 1 0.4 0.38 0.93 2.11 2.32 1.1
Full Finetuning 0.72 0.58 0.8 3.16 2.19 0.69 0.76 0.08 0.1 2.98 1.25 0.42
Prompt Tuning 0.86 +/- 0.0 0.75 +/- 0.01 0.87 +/- 0.01 3.18 +/- 0.02 2.44 +/- 0.02 0.77 +/- 0.01 0.82 +/- 0.04 0.86 +/- 0.01 1.05 +/- 0.06 2.93 +/- 0.02 1.98 +/- 0.1 0.67 +/- 0.04
LoRA 0.86 +/- 0.01 0.77 +/- 0.0 0.9 +/- 0.01 3.27 +/- 0.01 2.49 +/- 0.03 0.76 +/- 0.01 0.84 +/- 0.0 0.74 +/- 0.06 0.88 +/- 0.07 3.09 +/- 0.04 1.98 +/- 0.15 0.64 +/- 0.06
IA3 0.84 +/- 0.01 0.75 +/- 0.01 0.9 +/- 0.02 3.26 +/- 0.03 2.48 +/- 0.04 0.76 +/- 0.01 0.8 +/- 0.02 0.72 +/- 0.11 0.9 +/- 0.15 3.0 +/- 0.04 1.77 +/- 0.08 0.59 +/- 0.03
Most Frequent 0.75 - - 2.62 - - 0.73 - - 2.27 - -
(b) Results for the non-medical base model.
@@
No Finetuning 0.42 0.46 1.09 1.8 2.05 1.14 0.4 0.7 1.74 1.85 1.9 1.03
Full Finetuning 0.82 0.74 0.91 3.21 2.3 0.72 0.75 0.08 0.1 2.94 1.47 0.5
Prompt Tuning 0.82 +/- 0.02 0.72 +/- 0.05 0.87 +/- 0.04 3.14 +/- 0.03 2.3 +/- 0.03 0.73 +/- 0.0 0.73 +/- 0.05 0.5 +/- 0.22 0.69 +/- 0.31 2.86 +/- 0.02 1.67 +/- 0.17 0.58 +/- 0.06
LoRA 0.82 +/- 0.02 0.74 +/- 0.01 0.9 +/- 0.01 3.2 +/- 0.08 2.4 +/- 0.01 0.75 +/- 0.02 0.8 +/- 0.02 0.31 +/- 0.04 0.39 +/- 0.05 3.05 +/- 0.09 1.31 +/- 0.09 0.43 +/- 0.02
IA3 0.83 +/- 0.01 0.75 +/- 0.01 0.9 +/- 0.0 3.19 +/- 0.03 2.45 +/- 0.06 0.77 +/- 0.01 0.81 +/- 0.01 0.75 +/- 0.03 0.93 +/- 0.04 3.0 +/- 0.04 1.69 +/- 0.08 0.56 +/- 0.02
Most Frequent 0.75 - - 2.62 - - 0.73 - - 2.27 - -
(a) Results for the medical base model.
Body Part Shift Question Type Shift
OoD: Leg OoD: Organ System
Closed-Ended Open-Ended Closed-Ended Open-Ended
i.i.d. OoD RR i.i.d. OoD RR i.i.d. OoD RR i.i.d. OoD RR
No Finetuning 0.41 0.36 0.88 1.24 1.3 1.05 0.39 0.4 1.03 1.27 1.24 0.98
Prompt Tuning 0.75 +/- 0.01 0.7 +/- 0.01 0.94 +/- 0.01 2.66 +/- 0.1 2.09 +/- 0.02 0.79 +/- 0.02 0.67 +/- 0.03 0.44 +/- 0.01 0.66 +/- 0.02 2.32 +/- 0.05 1.39 +/- 0.07 0.6 +/- 0.04
LoRA 0.74 +/- 0.0 0.7 +/- 0.0 0.95 +/- 0.01 2.69 +/- 0.13 2.14 +/- 0.01 0.8 +/- 0.04 0.73 +/- 0.0 0.43 +/- 0.03 0.6 +/- 0.04 2.34 +/- 0.05 1.38 +/- 0.08 0.59 +/- 0.04
IA3 0.74 +/- 0.0 0.69 +/- 0.02 0.93 +/- 0.02 2.75 +/- 0.05 2.12 +/- 0.01 0.77 +/- 0.02 0.7 +/- 0.04 0.39 +/- 0.06 0.56 +/- 0.1 2.32 +/- 0.03 1.33 +/- 0.07 0.57 +/- 0.03
(b) Results for the non-medical base model.
Body Part Shift Question Type Shift
OoD: Leg OoD: Organ System
Closed-Ended Open-Ended Closed-Ended Open-Ended
i.i.d. OoD RR i.i.d. OoD RR i.i.d. OoD RR i.i.d. OoD RR
No Finetuning 0.2 0.19 0.95 1.21 1.15 0.95 0.19 0.35 1.83 1.14 1.33 1.16
Prompt Tuning 0.73 +/- 0.01 0.67 +/- 0.0 0.92 +/- 0.02 2.74 +/- 0.07 2.02 +/- 0.04 0.74 +/- 0.02 0.69 +/- 0.04 0.47 +/- 0.05 0.68 +/- 0.08 2.33 +/- 0.06 1.59 +/- 0.18 0.68 +/- 0.06
LoRA 0.73 +/- 0.03 0.68 +/- 0.04 0.93 +/- 0.02 2.69 +/- 0.14 2.12 +/- 0.01 0.79 +/- 0.04 0.73 +/- 0.0 0.43 +/- 0.01 0.6 +/- 0.02 2.32 +/- 0.03 1.35 +/- 0.04 0.58 +/- 0.03
IA3 0.72 +/- 0.03 0.66 +/- 0.03 0.93 +/- 0.02 2.63 +/- 0.12 2.08 +/- 0.03 0.79 +/- 0.03 0.72 +/- 0.02 0.36 +/- 0.08 0.5 +/- 0.11 2.36 +/- 0.01 1.3 +/- 0.1 0.55 +/- 0.04